OpenStudy (anonymous):

What is the specific heat of silver if it takes 66.8cal to raise the temperature of a 75g silver bar by 16.0∘C?

OpenStudy (shrutipande9):

Use d formula Q=mct 66.8=75*c*16 c=66*8/75*16

OpenStudy (shrutipande9):

66.8 I mean

OpenStudy (shrutipande9):

0.055cal/g deg C

OpenStudy (anonymous):

thank you very much

OpenStudy (shrutipande9):

Ur welcome:)